Any assistance is greatly appreciated in advance.
Q: Write a system of two equations in two variables to solve the problem.
A woman invested some money at 8% and some at 9% annual simple interest. The interest for 1 year on the combined investment of $10,000 was $880. How much was invested at each rate?
8% annual interest $ _____________

9% annual interest $ _____________

Answer by rothauserc(4718)

x + y = 10000
.08x +.09y = 880
solve first equation for x and substitute in second equation
.08(10000-y) +.09y = 880
800 -.08y +.09y = 880
.01y = 80
y = 8000
therefore,
$2000 was invested at 8% and $8000 was invested at 9%
check the answers
.08*2000 +.09*8000 = 880
160 + 720 = 880
880 = 880
answers check